OpenStudy (anonymous):

2,1,1/2,1/4 ..... A. arithmetic B. geometric C. both D. neither

OpenStudy (mathstudent55):

Do you know the difference between an arithmetic sequence and a geometric sequence?

OpenStudy (anonymous):

Yes .. is the answer Geometric ?

OpenStudy (mathstudent55):

Yes.

OpenStudy (anonymous):

Alright thanks :D

OpenStudy (mathstudent55):

1/2 = 1/2 (1/2) / 1 = 1/2 (1/4) / (1/2) = 1/2 Every term divided by the previous term gives the same result. That is the common ratio. That means it's a geometric sequence.

OpenStudy (mathstudent55):

You're welcome.